2- Choose the correct answer (Write down the correct answer letter AND valuc at your answer booklet) a. Let X be a real-valued, continuous random varia ble and let YX. Then, If y 2 0, then the cumulative distribution function E,(v): F,(一./5) _ Fx(,5) (D) -Fx(JF) _ F,(-(y) (E) none of the above pt. If a random variable X has a PDF (x)-2(x-1) 1ss2 (A) 35 (B) 5/3 variance of 75. The probability that X lies between 40 and 50 is: 0 elsewhere then E(x) is: (C) 4/5 (D) 5/4(E) none of the above 2pt. 0 elsewhere c. If the continuous random variable X is uniformly distributed with a mean of 45 and a (C) 2/3 (D) 12 1ss2 0 elsewhere (E) none of the above 2pt. , then ơ1 is: If a random variable X has a PDFx)-1) 0 (A) 1/20 (B) 20 (C) 1/18 (D) 1/24(E) Done of the above e | If random variable X has a mean 5 , a standard deviation ơ,-4, and x 2-2X1hen: (A) ,-,-3 and ơ rt: 4 (D)-w. .3 and ơ r-5 (E) none of theabove. . .2pt.

Choose the correct answer with detailed explanation please .
Answer #1

Here' the answer to the question. Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

2.

a. Y=X^2
So, the cumulative distribution lies between +/- X or +/- sqrt(Y)
So, Ey(y) = Fx(x) - Fx(-x) = Fxsqrt(y) - Fx(-sqrt(y))
So, B is correct


b. E(x) = intergral of x*f(x) limits 1 to 2 = integral of x*(2*(x-1))
= integration of 2x^2 - 2x
2x^3/3 - 2x^2/2 = 2x^3/3 -x^2 with limits 1 to 2:

(2*(2^3)/3 - 2^2)-(2*1^3/3 -1^2 )
= (16/3 - 4)- (-1/3)
= (16-12)/3 + 1/3
= 5/3

B is correct

c. This is basically p(X) = (x-a)/(b-a) = ?
Given : Mean = (a+b)/2 = 45: So, (a+b) = 90
variance = 75= (b-a)^2 / 12
(b-a)^2 = 75*12
(b-a) = 5*6 = 30

Adding (a+b) = 90, (b-a) = 30, 2b = 120, b = 60, a= 30
P(40<X<50) = (x-a)/(b-a) = (50-40)/(60-30) = 10/30 = 1/3
A is correct

d. This is Var(x) = E(X^2) - E(X)^2
= Integral of x^2f(x) - (5/3)^2 ( from part b calculated above)
= Integral of x^2*(2*(x-1)) -(5/3)^2
= Integral of 2x^3 -2*x^2 = x^4/2 - 2/3 *x^3 from 1 to 2 limits - (5/3)^2
((2^4)/2 - (2/3)*(2^3) )-((1^4)/2 - (2/3)*(1^3)) - (25/9)
= (8-16/3) - (1/2 - (2/3)*(1)) - (25/9)
= 8/3 - (1/2) + (2/3) - (25/9)
= (48-9+12-50)/(18)
= 1/18

C is correct

e. Y = 2-2X
Mean of Y = E(Y) = E(2-2X) = E(2) + E(-2X) = 2 - 2*E(X) = 2-2*(5) = -8

Expected value of constant is same constant. So, E(2) = 2, and E(aX) = a*E(x)

Stdev(Y) = Stdev(2-2X) = Stdev(2) + Stdev(2x) = 0+ 2*Stdev(X) = +2*4 = 8
Expected value of constant is 0. So, Stdev(2) = 0, and Stdev(aX) = a*Stdev(x)

B is correct